• Main INDEX
  • Monthly INDEX
  • PREV
  • NEXT

    User name adaq

    Log entry time 08:03:06 on April 19, 2004

    Entry number 118566

    keyword=Owl Shift Summary

    Owl shift for Monday April 19th, 2004


    Leader:   B. Moffit
    Target:   J. Singh
    Worker:  S. Choi

    Summary

      Smooth running most of the night. Lost some time due to beam trips, RICH HV trip, and a slight change to the Left Septum current.

    Significant Events

    • 00:00 Began the shift with the end of Run 2542-Production Run.
    • 00:01 Began Run 2543-Production Run
    • 00:52 Began Run 2544-Production Run
    • 01:59 Began Run 2545-Production Run
    • 02:57 Began Run 2546-Production Run
    • 03:50 Began Run 2547-Production Run
    • 04:45 Increased Left Septum current slightly to make it agree with current used before the septum tripped yesterday.
      See halog entries (118545 and 118548)
    • 04:55 Began Run 2548-Production Run
    • 05:16 RICH HV Trip. See halog entry 118549
    • 05:22 Began Run 2549-Production Run
    • 06:12 Began Run 2550-Production Run
    • 06:36 Asked MCC for 75uA.
    • 06:46 Began Run 2551-Production Run
    • 06:50 Asked MCC for 100uA.
    • 06:51 Began Run 2552-Production Run
    • 07:00 MCC: No more CW. Tune beam into hall to check optics.
    • 07:20 Seonho Leaves.

    Runs Taken

    • Run 2543 - Production Run (12C)
    • Run 2544 - Production Run (12C)
    • Run 2545 - Production Run (12C)
    • Run 2546 - Production Run (12C)
    • Run 2547 - Production Run (12C)
    • Run 2548 - Production Run (12C)
    • Run 2549 - Production Run (12C)
    • Run 2550 - Production Run (12C)
    • Run 2551 - Production Run (12C)
    • Run 2552 - Production Run (12C)

    Halog Ratios:
    a-onl  : 21/39 = 53.8%
    Seonho :  1/39 =  2.6%
    Bryan  :  3/39 =  7.7%
    Jaideep: 14/39 = 35.9%